# Electromagnetic Multipole Theory for Two-dimensional Photonics

### v1

@@ Abstract
The multipole analysis has been widely employed to investigate a broad spectrum of cutting-edge phenomena in three-dimensional (3-D) structures, including meta-atoms for the manipulation of EM waves 1, design of metadimers with unique optical properties 2, Kerker scattering to control the directionality of light via multipolar interferences 3, zero optical back-scattering from single nanoparticles 4, directed light emission thru multipolar interferences 5, coupling enhancement of multipole resonances via optical beams 6, generalized Kerker effects in meta-optics 7, selective excitation of multipolar resonances via cylindrical vector beams 8, magnetic switching of Kerker scattering 9, dynamic control for invisibility to superscattering switching 10, excitation of higher-order multipolar modes via displacement resonance 11, or more complex 3-D structures consisting of meta-atoms or meta-molecules such as metasurfaces and metagratings 12, 13, 14, 15, 16.
@@
On the other hand, two-dimensional (2-D) structures, such as nanowires, long rods, oligomers, or metalattices/metagratings thereof 17, exhibit unique optical properties that can be harnessed for advanced technological applications. The multipole analysis of cylindrical structures has a plethora of applications, including efficient magnetic mirrors 18, scattering invisibility in all-dielectric nanoparticles 19, active tuning of directional scattering of magneto-optical structures 20, polarization manipulation in cylindrical metalattices 21, optical beam steering control via dielectric diffractionless metasurfaces and diffractive metagratings 22, tuning of toroidal dipole resonances in all-dielectric nanocylindrical metamolecules 23, reconfigurable metalattices via magneto-optically coated rods 24, arbitrary beam steering exploiting transverse Kerker scattering 25, and extreme nonreciprocal scattering in asymmetric gyrotropic cylindrical trimers 26.
@@
Complete theories for the multipole decomposition have been developed exclusively for 3-D particles of isotropic materials. These theories include the expansion of the EM field in spherical eigenvectors 27, exact expressions for source dipolar moments as well as for dipoles that radiate a definite polarization handedness 28, the derivation of exact multipole moments beyond the subwavelength limit 29, and an efficient multipole decomposition procedure using Lebedev and Gaussian quadrature 30.
@@
In this work, we provide, for the first time, a full-wave theoretical framework for the derivation of multipole decomposition in arbitrarily 2-D inhomogeneous gyrotropic structures composed either of an isolated scatterer or collections of such. At first, the scattered electric field ${\mathbf{E}}^{\rm sc}({\boldsymbol{\rho}})$—where ${\boldsymbol{\rho}}=(\rho,\varphi)$ is the position vector in polar coordinates—is expanded in divergenceless cylindrical vector wave functions (CVWF) with unknown expansion coefficients $A_{m}$, $B_{m}$ that determine the multipole response. $A_{m}$, $B_{m}$ are next expressed via contour integrals with integrand functions considering the vectorial components of ${\mathbf{E}}^{\rm sc}({\boldsymbol{\rho}})$. In sequence, ${\mathbf{E}}^{\rm sc}({\boldsymbol{\rho}})$ is calculated through a 2-D electric field volume integral equation (EFVIE) with kernels the products of the tensorial 2-D Green’s function (GF) expansion ${\mathbb{G}}({\boldsymbol{\rho}}-{\boldsymbol{\rho}}^{\prime})$ and the equivalent 2-D volumetric electric ${\mathbf{J}}_{\rm eq}({\boldsymbol{\rho}})$ and magnetic ${\mathbf{M}}_{\rm eq}({\boldsymbol{\rho}})$ current densities. This scheme allows for the determination of $A_{m}$ and $B_{m}$ via 2-D volumetric integrations in the gyrotropic region of the scatterers.
@@
We exhaustively validate our theory, first by calculating the multipole decomposition of various 2-D structures whose scattered field has been obtained from the full-wave Maxwell-solver COMSOL. We also compare our results to those obtained with the exact formulation for isotropic and gyrotropic circular cylinders 31, with Mathieu functions method for elliptical cylinders 32, with the coupled-field volume integral equation-cylindrical Dini series expansion (CFVIE-CDSE) method for core-shell circular cylinders 33, and with the method of auxiliary sources (MAS) for circular core-elliptical shell and circular dimer cylinders 34. In all these comparisons, an excellent agreement is observed.
@@
The paper is organized as follows: Section 1 develops the theoretical framework, Section 2 provides various validation examples, Section 3 discusses a photonics application on oligomers-based highly directional switching using active media, and Section 4 concludes the paper.
@@ 1.1 Multipole Decomposition for an Isolated Scatterer
The problem configuration is depicted in Figure 1(a). An incident electric field ${\mathbf{E}}^{\rm inc}({\boldsymbol{\rho}})$ is impinging on a 2-D isolated scatterer $S$ described by an inhomogeneous, gyrotropic material with constitutive parameters
@@
where $A_{m}$, $B_{m}$ are the unknown expansion coefficients to be evaluated, $Z_{0}=(\mu_{0}/\epsilon_{0})^{1/2}$ is free space impedance, $k_{0}=\omega(\epsilon_{0}\mu_{0})^{1/2}$ is free space wavenumber, and ${\mathbf{M}}^{(4)}_{m}(k_{0},{\boldsymbol{\rho}})$, ${\mathbf{N}}^{(4)}_{m}(k_{0},{\boldsymbol{\rho}})$ are the CVWF of the fourth kind which, for infinitely long configurations—i.e., $\partial/\partial z=0$—are given by 33, 36
@@
In eq 1.1, $H_{m}$ is the Hankel function of the second kind—the superscript $(2)$ is omitted for simplicity—and $H^{\prime}_{m}$ is the derivative of $H_{m}$ with respect to its argument. On the right-hand side of eq 2, the term that involves the ${\mathbf{M}}^{(4)}_{m}(k_{0},{\boldsymbol{\rho}})$ CVWF represents the TE solution while the term that involves the ${\mathbf{N}}^{(4)}_{m}(k_{0},{\boldsymbol{\rho}})$ CVWF represents the TM solution. For TE scattering, $m=0$ gives the magnetic dipolar (MD) response, $m=\pm 1$ the electric dipolar (ED) response, and $m=\pm 2$ the electric quadrupolar (EQ) response; for TM scattering, the ED, MD, and MQ contributions are given by $m=0$, $m=\pm 1$ and $m=\pm 2$, respectively 19. We aim to calculate $A_{m}$ and $B_{m}$ to determine the multipole response fully.
@@
As evident, $B_{m}$ is evaluated either via $E^{\rm sc}_{\rho}$ or $E^{\rm sc}_{\varphi}$. To evaluate the components of ${\mathbf{E}}^{\rm sc}({\boldsymbol{\rho}})$, the 2-D scattering problem is formulated by the EFVIE 33, namely
@@
In eq 1.1, $T$ denotes transposition, $g({\boldsymbol{\rho}}-{\boldsymbol{\rho}}^{\prime})=-i/4H_{0}(k_{0}|{\boldsymbol{\rho}}-{\boldsymbol{\rho}}^{\prime}|)$ is the 2-D free space GF 37, and ${\mathbb{X}}_{\rm e}({\boldsymbol{\rho}})=\boldsymbol{\epsilon}({\boldsymbol{\rho}})/\epsilon_{0}-{\mathbb{I}}$ and ${\mathbb{X}}_{\rm m}({\boldsymbol{\rho}})=\boldsymbol{\mu}({\boldsymbol{\rho}})/\mu_{0}-{\mathbb{I}}$ the normalized tensorial electric and magnetic contrast function. When ${\boldsymbol{\rho}}\in S_{0}$, the two 2-D volumetric integrals in eq 1.1 represent the ${\mathbf{E}}^{\rm sc}({\boldsymbol{\rho}})$; introducing the 2-D volumetric current densities ${\mathbf{J}}_{\rm eq}({\boldsymbol{\rho}})$ and ${\mathbf{M}}_{\rm eq}({\boldsymbol{\rho}})$, one writes
@@
To evaluate ${\mathbf{E}}^{\rm sc}_{J}({\boldsymbol{\rho}})$, the tensorial expansion ${\mathbb{G}}({\boldsymbol{\rho}}-{\boldsymbol{\rho}}^{\prime})$ of the 2-D GF $g({\boldsymbol{\rho}}-{\boldsymbol{\rho}}^{\prime})$ is used, given by eq 6 of ref 33. Since the 2-D volumetric integrals in eq 1.1 are evaluated for ${\boldsymbol{\rho}}^{\prime}\in S$, and because ${\boldsymbol{\rho}}\in S_{0}$, the branch $\rho>\rho^{\prime}$ is used for the ${\mathbb{G}}({\boldsymbol{\rho}}-{\boldsymbol{\rho}}^{\prime})$. This branch has the tensorial expansion
@@ 1.2 Multipole Decomposition for a Collection of Scatterers
In this case, the configuration comprises $N$ scatterers located in free space, as depicted in Figure 2. Defining a global polar coordinate system ${\rm O}xy$, each scatterer occupies domain $S_{j}$, $j=1,2,\ldots,N$, and it is characterized by a gyrotropic material having $\boldsymbol{\epsilon}_{j}({\boldsymbol{\rho}})$, $\boldsymbol{\mu}_{j}({\boldsymbol{\rho}})$, $j=1,2,\ldots,N$. The background medium $S_{0}$ is free space. Upon external excitation , the 2-D volumetric current densities ${\mathbf{J}}_{{\rm eq},j}({\boldsymbol{\rho}})$, ${\mathbf{M}}_{{\rm eq},j}({\boldsymbol{\rho}})$, ${\boldsymbol{\rho}}\in S_{j}$, $j=1,2,\ldots,N$, are induced. The collective response of all equivalent current densities produces the scattered field ${\mathbf{E}}^{\rm sc}({\boldsymbol{\rho}})$ in $S_{0}$, i.e.,
@@
1.3 Scattering Width and Total Scattering Cross Section
@@
In the far-field, ${\mathbf{E}}^{\rm sc}({\boldsymbol{\rho}})\sim E_{0}e^{-ik_{0}\rho}/\sqrt{\rho}{\mathbf{f}}(\varphi)$, $\rho\rightarrow\infty$, where ${\mathbf{f}}(\varphi)$ is the scattering amplitude. Then, the scattering width $\sigma(\varphi)$, $0\leqslant\varphi<2\pi$, is given by 35
@@
The full-wave total scattering cross-section is given by
@@
where $\sigma_{\rm d}(\varphi)=|1/\sqrt{k_{0}}\tilde{f}_{\varphi}(\varphi)|^{2}+|1/\sqrt{k_{0}}\tilde{f}_{z}(\varphi)|^{2}$ is the differential scattering cross-section. The $Q_{\rm sc}$ for TE scattering is calculated solely via $\tilde{B}_{m}$ while for TM scattering via $\tilde{A}_{m}$. For TE scattering, the $m=0$ term in eq 24 gives the MD response, the $m=\pm 1$ terms the ED response, and the $m=\pm 2$ terms the EQ response to the spectrum. For TM scattering, the ED, MD, and MQ responses to the spectrum are calculated by keeping the $m=0$, $m=\pm 1$, and $m=\pm 2$ terms, respectively, in eq 24.
@@ 2 Validation
Herein, we validate the developed theory by calculating the multipole decomposition of various 2-D structures and compare the results with alternative formulations. In particular, the present theory is combined with COMSOL that we use to compute the full-wave $Q_{\rm sc}$ spectrum vs. the operating frequency $f$ via eq 23, and its multipole contributions using eq 24, eq 19, eq 16, eq 15, eq 1.1, eq 1.1, and eq 9. Then, we compare our results with the exact formulation for isotropic and gyrotropic circular cylinders 31, with Mathieu functions method for elliptical cylinders 32, with the CFVIE-CDSE for core-shell circular cylinders 33, and with the MAS for circular core-elliptical shell and circular dimer cylinders 34.
@@
At first, we focus on the visible and near-IR parts of the spectrum. In Figure 3(a), we assume a high-index dielectric cylinder of circular cross-section. The values of the parameters are given in the figure caption. We calculate the normalized full-wave total scattering cross-section, i.e., $Q_{\rm sc}/a$, using the exact solution 31, under TE illumination, as shown in Figure 3(b) by the black curve. The multipole decomposition is inherent in the analytical formulation, and the red and green curves show the corresponding MD and ED contributions, respectively. The calculated full-wave $Q_{\rm sc}/a$ from eq 23—with the aid of COMSOL—is depicted with black dots, while the corresponding multipole decomposition using the methodology proposed here is depicted by red—MD—and green—ED—dots. The agreement between the exact calculation and the multipole decomposition of the developed theory is apparent. The agreement is also true for TM illumination, as shown in Figure 3(c).
@@
Next, we assume a high-index dielectric cylinder with an elliptical cross-section, as shown in Figure 4(a). The semi-major and semi-minor axes are $a$ and $b$, respectively. We keep $a=20~{}\mu{\rm m}$ and consider two different values for $b$, i.e., $b=0.6a$ and $b=0.7a$. We assume TE illumination for both cases. In Figure 4(b), we depict the $Q_{\rm sc}/a$ spectrum when $b=0.6a$. The black curve corresponds to the full-wave $Q_{\rm sc}/a$ using Mathieu functions method 32, that can be decomposed to its MD—red curve—and ED—green curve—contributions. The corresponding full-wave $Q_{\rm sc}/a$ via eq 23 is shown by black dots, while the result of the proposed multipole decomposition is shown by red—MD—and green—ED—dots. The proposed theory matches perfectly that from 32. In Figure 4(c), we show the respective results for the system with $b=0.7a$.
@@
Figure 6: (a) TE scattering by a core-shell circular cylinder. Values of parameters: $\epsilon_{cr}=25$, $\mu_{cr}=1$, $\epsilon_{1r}=4$, $\epsilon_{2r}=1$, $\epsilon_{3r}=5$, $\mu_{1r}=2$, $\mu_{2r}=0.5$, $\mu_{3r}=3$, $a=20~{}\mu{\rm m}$, $b/a=0.75$. (b) $Q_{\rm sc}/a$ spectrum. Curves: CFVIE-CDSE; dots: this work via COMSOL; black: full-wave; red: MD; green: ED.
Next, we focus our study on the THz-spectrum of frequencies. We proceed with an inhomogeneous core-shell cylinder of the circular cross-section shown in Figure 6(a), consisting of a high-index dielectric $b$-radius core of relative permittivity $\epsilon_{cr}$ and relative permeability $\mu_{cr}$, coated with a gyrotropic shell of outer radius $a$. The full-wave $Q_{\rm sc}/a$ for the TE illumination is calculated using the CFVIE-CDSE 33. Results are shown by the black curve in Figure 6(b). In contrast, the colored curves depict the corresponding MD—red—and ED—green—contributions. The multipole decomposition is shown by dots using the respective colors. It is evident that all results agree perfectly.

### v2

@@ Abstract
The multipole analysis has been widely employed to investigate a broad spectrum of cutting-edge phenomena in three-dimensional (3-D) structures, including meta-atoms for the manipulation of EM waves 1, design of metadimers with unique optical properties 2, Kerker scattering to control the directionality of light via multipolar interferences 3, zero optical back-scattering from single nanoparticles 4, directed light emission thru multipolar interferences 5, coupling enhancement of multipole resonances via optical beams 6, highly transmissive metasurfaces for polarization control7, generalized Kerker effects in meta-optics 8, selective excitation of multipolar resonances via cylindrical vector beams 9, magnetic switching of Kerker scattering 10, dynamic control for invisibility to superscattering switching 11, excitation of higher-order multipolar modes via displacement resonance 12, or more complex 3-D structures consisting of meta-atoms or meta-molecules such as metasurfaces and metagratings 13, 14, 15, 16, 17, as well as dense ensembles of plasmonic nanoparticles18.
@@
On the other hand, two-dimensional (2-D) structures, such as nanowires, long rods, oligomers, or metalattices/metagratings thereof 19, exhibit unique optical properties that can be harnessed for advanced technological applications. The multipole analysis of cylindrical structures has a plethora of applications, including efficient magnetic mirrors 20, scattering invisibility in all-dielectric nanoparticles 21, active tuning of directional scattering of magneto-optical structures 22, polarization manipulation in cylindrical metalattices 23, optical beam steering control via dielectric diffractionless metasurfaces and diffractive metagratings 24, tuning of toroidal dipole resonances in all-dielectric nanocylindrical metamolecules 25, reconfigurable metalattices via magneto-optically coated rods 26, arbitrary beam steering exploiting transverse Kerker scattering 27, and extreme nonreciprocal scattering in asymmetric gyrotropic cylindrical trimers 28.
@@
Complete theories for the multipole decomposition have been developed exclusively for 3-D particles of isotropic materials. These theories include the expansion of the EM field in spherical eigenvectors 29, exact expressions for source dipolar moments as well as for dipoles that radiate a definite polarization handedness 30, the derivation of exact multipole moments beyond the subwavelength limit 31, an efficient multipole decomposition procedure using Lebedev and Gaussian quadrature 32, and exact multipole moments for magnetic particles of arbitrary shape and size33.
@@
In this work, we provide, for the first time, a full-wave theoretical framework for the derivation of multipole decomposition in arbitrarily 2-D inhomogeneous gyrotropic structures composed either of an isolated scatterer or collections of such. This framework is undoubtedly significant for various fields of photonics, including nanoparticle engineering34, dielectric metasurfaces35, superscattering enhancement11 and Mie-tronics36, to name a few. In addition, being an extensive theory, it complements existing theories on multiple scattering by nanoparticle structures37, 38 applied for photonics applications. At first, the scattered electric field ${\mathbf{E}}^{\rm sc}({\boldsymbol{\rho}})$—where ${\boldsymbol{\rho}}=(\rho,\varphi)$ is the position vector in polar coordinates—is expanded in divergenceless cylindrical vector wave functions (CVWF) with unknown expansion coefficients $A_{m}$, $B_{m}$ that determine the multipole response. $A_{m}$, $B_{m}$ are next expressed via contour integrals with integrand functions considering the vectorial components of ${\mathbf{E}}^{\rm sc}({\boldsymbol{\rho}})$. In sequence, ${\mathbf{E}}^{\rm sc}({\boldsymbol{\rho}})$ is calculated through a 2-D electric field volume integral equation (EFVIE) with kernels the products of the tensorial 2-D Green’s function (GF) expansion ${\mathbb{G}}({\boldsymbol{\rho}}-{\boldsymbol{\rho}}^{\prime})$ and the equivalent 2-D volumetric electric ${\mathbf{J}}_{\rm eq}({\boldsymbol{\rho}})$ and magnetic ${\mathbf{M}}_{\rm eq}({\boldsymbol{\rho}})$ current densities. This scheme allows for the determination of $A_{m}$ and $B_{m}$ via 2-D volumetric integrations in the gyrotropic region of the scatterers.
@@
We exhaustively validate our theory, first by calculating the multipole decomposition of various 2-D structures whose scattered field has been obtained from the full-wave Maxwell-solver COMSOL. We also compare our results to those obtained with the exact formulation for isotropic and gyrotropic circular cylinders 39, with Mathieu functions method for elliptical cylinders 40, with the coupled-field volume integral equation-cylindrical Dini series expansion (CFVIE-CDSE) method for core-shell circular cylinders 41, and with the method of auxiliary sources (MAS) for circular core-elliptical shell and circular dimer cylinders 42. In all these comparisons, an excellent agreement is observed. The main advantage that our approach provides, as compared to the above mentioned ones39, 40, 41, 42 used for method verification, is that it allows the theory to be implemented in any general purpose solver, such as COMSOL, so as to perform the multipole decomposition on complicated structures composed of scatterers of non-canonical shapes with anisotropic material properties. This is a burdensome task for semi-analytical techniques, since their development requires cumbersome manipulations, while their applicability is limited, mainly, to canonical shapes, or perturbed variants.
@@
The paper is organized as follows: In the first section, we develop the theoretical framework, then we provide various validation examples, after which there is a discussion on a photonics application on oligomers-based highly directional switching using active media, while the final section concludes the paper.
@@ 1.1 Multipole Decomposition for an Isolated Scatterer
The configuration problem is depicted in Figure 1(a). An incident electric field ${\mathbf{E}}^{\rm inc}({\boldsymbol{\rho}})$ is impinging on a 2-D isolated scatterer $S$ described by an inhomogeneous, gyrotropic material with constitutive parameters
@@
where $A_{m}$, $B_{m}$ are the unknown expansion coefficients to be evaluated, $Z_{0}=(\mu_{0}/\epsilon_{0})^{1/2}$ is free space impedance, $k_{0}=\omega(\epsilon_{0}\mu_{0})^{1/2}$ is free space wavenumber, and ${\mathbf{M}}^{(4)}_{m}(k_{0},{\boldsymbol{\rho}})$, ${\mathbf{N}}^{(4)}_{m}(k_{0},{\boldsymbol{\rho}})$ are the CVWF of the fourth kind which, for infinitely long configurations—i.e., $\partial/\partial z=0$—are given by 41, 44
@@
In eq 1.1, $H_{m}$ is the Hankel function of the second kind—the superscript $(2)$ is omitted for simplicity—and $H^{\prime}_{m}$ is the derivative of $H_{m}$ with respect to its argument. On the right-hand side of eq 2, the term that involves the ${\mathbf{M}}^{(4)}_{m}(k_{0},{\boldsymbol{\rho}})$ CVWF represents the TE solution, while the term that involves the ${\mathbf{N}}^{(4)}_{m}(k_{0},{\boldsymbol{\rho}})$ CVWF represents the TM solution. For TE scattering, $m=0$ gives the magnetic dipolar (MD) response, $m=\pm 1$ the electric dipolar (ED) response, and $m=\pm 2$ the electric quadrupolar (EQ) response; for TM scattering, the ED, MD, and MQ contributions are given by $m=0$, $m=\pm 1$ and $m=\pm 2$, respectively 21. We aim to calculate $A_{m}$ and $B_{m}$ to determine the multipole response fully.
@@
As evident, $B_{m}$ is evaluated either via $E^{\rm sc}_{\rho}$ or $E^{\rm sc}_{\varphi}$. To evaluate the components of ${\mathbf{E}}^{\rm sc}({\boldsymbol{\rho}})$, the 2-D scattering problem is formulated by the EFVIE 41, namely
@@
In eq 1.1, $T$ denotes transposition, $g({\boldsymbol{\rho}}-{\boldsymbol{\rho}}^{\prime})=-i/4H_{0}(k_{0}|{\boldsymbol{\rho}}-{\boldsymbol{\rho}}^{\prime}|)$ is the 2-D free space GF 45, and ${\mathbb{X}}_{\rm e}({\boldsymbol{\rho}})=\boldsymbol{\epsilon}({\boldsymbol{\rho}})/\epsilon_{0}-{\mathbb{I}}$ and ${\mathbb{X}}_{\rm m}({\boldsymbol{\rho}})=\boldsymbol{\mu}({\boldsymbol{\rho}})/\mu_{0}-{\mathbb{I}}$ the normalized tensorial electric and magnetic contrast function. When ${\boldsymbol{\rho}}\in S_{0}$, the two 2-D volumetric integrals in eq 1.1 represent the ${\mathbf{E}}^{\rm sc}({\boldsymbol{\rho}})$; introducing the 2-D volumetric current densities ${\mathbf{J}}_{\rm eq}({\boldsymbol{\rho}})$ and ${\mathbf{M}}_{\rm eq}({\boldsymbol{\rho}})$, one writes
@@
To evaluate ${\mathbf{E}}^{\rm sc}_{J}({\boldsymbol{\rho}})$, the tensorial expansion ${\mathbb{G}}({\boldsymbol{\rho}}-{\boldsymbol{\rho}}^{\prime})$ of the 2-D GF $g({\boldsymbol{\rho}}-{\boldsymbol{\rho}}^{\prime})$ is used, given by eq 6 of ref 41. Since the 2-D volumetric integrals in eq 1.1 are evaluated for ${\boldsymbol{\rho}}^{\prime}\in S$, and because ${\boldsymbol{\rho}}\in S_{0}$, the branch $\rho>\rho^{\prime}$ is used for the ${\mathbb{G}}({\boldsymbol{\rho}}-{\boldsymbol{\rho}}^{\prime})$. This branch has the tensorial expansion
@@ 1.2 Multipole Decomposition for a Collection of Scatterers
In this case, the configuration comprises $N$ scatterers located in free space, as depicted in Figure 2. Defining a global polar coordinate system ${\rm O}xy$, each scatterer occupies domain $S_{j}$, $j=1,2,\ldots,N$, and is characterized by a gyrotropic material having $\boldsymbol{\epsilon}_{j}({\boldsymbol{\rho}})$, $\boldsymbol{\mu}_{j}({\boldsymbol{\rho}})$, $j=1,2,\ldots,N$. The background medium $S_{0}$ is free space. Upon external excitation , the 2-D volumetric current densities ${\mathbf{J}}_{{\rm eq},j}({\boldsymbol{\rho}})$, ${\mathbf{M}}_{{\rm eq},j}({\boldsymbol{\rho}})$, ${\boldsymbol{\rho}}\in S_{j}$, $j=1,2,\ldots,N$, are induced. The collective response of all equivalent current densities produces the scattered field ${\mathbf{E}}^{\rm sc}({\boldsymbol{\rho}})$ in $S_{0}$, i.e.,
@@
1.3 Scattering Width, Scattering Cross-Section, Absorption Cross-Section, and Extinction Cross-Section
@@
In the far-field, ${\mathbf{E}}^{\rm sc}({\boldsymbol{\rho}})\sim E_{0}e^{-ik_{0}\rho}/\sqrt{\rho}{\mathbf{f}}(\varphi)$, $\rho\rightarrow\infty$, where ${\mathbf{f}}(\varphi)$ is the scattering amplitude. Then, the scattering width $\sigma(\varphi)$, $0\leqslant\varphi<2\pi$, is given by 43
@@
The full-wave scattering cross-section is given by
@@
where $\sigma_{\rm d}(\varphi)=|1/\sqrt{k_{0}}\tilde{f}_{\varphi}(\varphi)|^{2}+|1/\sqrt{k_{0}}\tilde{f}_{z}(\varphi)|^{2}$ is the differential scattering cross-section. The $Q_{\rm sc}$ for TE scattering is calculated solely via $\tilde{B}_{m}$, while for TM scattering, via $\tilde{A}_{m}$. For TE scattering, the $m=0$ term in eq 24 gives the MD response, the $m=\pm 1$ terms the ED response, and the $m=\pm 2$ terms the EQ response to the spectrum. For TM scattering, the ED, MD, and MQ responses to the spectrum are calculated by keeping the $m=0$, $m=\pm 1$, and $m=\pm 2$ terms, respectively, in eq 24.
Finally, to discuss the specifics of the scattering from the point of view of energy relations, we introduce the absorption cross-section $Q_{\rm abs}$ whose full-wave evaluation is given by
$Q_{\rm abs}=\frac{1}{|{\mathbf{S}}^{\rm inc}({\boldsymbol{\rho}})|}\sum_{j=1}^{N}\int_{{\boldsymbol{\rho}}^{\prime}\in S_{j}}Q_{j}{\rm d}{\boldsymbol{\rho}}^{\prime},$ (25)
where $Q_{j}$ is the power loss density of each domain $S_{j}$. Therefore, the full-wave extinction cross-section is given by $Q_{\rm ext}=Q_{\rm sc}+Q_{\rm abs}$46, while its multipole decomposition is obtained via46
$Q_{\rm ext}=\frac{4}{k_{0}}\sum_{m=-\infty}^{\infty}{\rm Re}\big{\{}\tilde{B}_{m}+\tilde{A}_{m}\big{\}},$ (26)
where ${\rm Re}$ denotes the real part.
@@ 2 Validation
Herein, we validate the developed theory by calculating the multipole decomposition of various 2-D structures and compare the results with alternative formulations. In particular, the present theory is combined with COMSOL which we use to compute the full-wave $Q_{\rm sc}$ spectrum vs. the operating frequency $f$ via eq 23, and its multipole contributions using eq 24, eq 19, eq 16, eq 15, eq 1.1, eq 1.1, and eq 9. Then, we compare our results with the exact formulation for isotropic and gyrotropic circular cylinders 39, with Mathieu functions method for elliptical cylinders 40, with the CFVIE-CDSE for core-shell circular cylinders 41, and with the MAS for circular core-elliptical shell and circular dimer cylinders 42.
@@
At first, we focus on the visible and near-IR parts of the spectrum. In Figure 3(a), we assume a high-index dielectric cylinder of circular cross-section. The values of the parameters are given in the figure caption. We calculate the normalized full-wave scattering cross-section, i.e., $Q_{\rm sc}/a$, using the exact solution 39, under TE illumination, as shown in Figure 3(b) by the black curve. The multipole decomposition is inherent in the analytical formulation, and the red and green curves show the corresponding MD and ED contributions, respectively. The calculated full-wave $Q_{\rm sc}/a$ from eq 23—with the aid of COMSOL—is depicted with black dots, while the corresponding multipole decomposition, using the methodology proposed here, is depicted by red—MD—and green—ED—dots. The agreement between the exact calculation and the multipole decomposition of the developed theory is apparent. The agreement is also true for TM illumination, as shown in Figure 3(c).
@@
Next, we assume a high-index dielectric cylinder with an elliptical cross-section, as shown in Figure 4(a). The semi-major and semi-minor axes are $a$ and $b$, respectively. We keep $a=20~{}\mu{\rm m}$ and consider two different values for $b$, i.e., $b=0.6a$ and $b=0.7a$. We assume TE illumination for both cases. In Figure 4(b), we depict the $Q_{\rm sc}/a$ spectrum when $b=0.6a$. The black curve corresponds to the full-wave $Q_{\rm sc}/a$ using Mathieu functions method 40, that can be decomposed to its MD—red curve—and ED—green curve—contributions. The corresponding full-wave $Q_{\rm sc}/a$ via eq 23 is shown by black dots, while the result of the proposed multipole decomposition is shown by red—MD—and green—ED—dots. The proposed theory matches perfectly the one from 40. In Figure 4(c), we show the respective results for the system with $b=0.7a$.
@@
Figure 6: (a) TE scattering by a core-shell circular cylinder. Values of parameters: $\epsilon_{cr}=25$, $\mu_{cr}=1$, $\epsilon_{1r}=4$, $\epsilon_{2r}=1$, $\epsilon_{3r}=5$, $\mu_{1r}=2$, $\mu_{2r}=0.5$, $\mu_{3r}=3$, $a=20~{}\mu{\rm m}$, $b/a=0.75$. (b) $Q_{\rm sc}/a$ spectrum. Curves: CFVIE-CDSE; dots: this work via COMSOL; black: full-wave; red: MD; green: ED. Orange dotted arrow: location of the first MD-ED intersection at $f=1.0389$ THz. (c) $\sigma/a$ at $f=1.0389$ THz. Orange curve: CFVIE-CDSE; orange dots: this work via COMSOL.
Next, we focus our study on the THz-spectrum of frequencies. We proceed with an inhomogeneous core-shell cylinder of the circular cross-section shown in Figure 6(a), consisting of a high-index dielectric $b$-radius core of relative permittivity $\epsilon_{cr}$ and relative permeability $\mu_{cr}$, coated with a gyrotropic shell of outer radius $a$. The full-wave $Q_{\rm sc}/a$ for the TE illumination is calculated using the CFVIE-CDSE 41. Results are shown by the black curve in Figure 6(b). In contrast, the colored curves depict the corresponding MD—red—and ED—green—contributions. The multipole decomposition is shown by dots using the respective colors. It is evident that all results agree perfectly. In addition, in Figure 6(c), we demonstrate the Kerker effect33, i.e., zero back-scattering (first Kerker point47), by plotting the full-wave normalized scattering width, i.e., $\sigma/a$, as computed by the CFVIE-CDSE and COMSOL. The Kerker effect takes place at $f=1.0389$ THz, as shown by the orange dotted arrow in Figure 6(b) where the MD-ED contributions intersect. Both methods in Figure 6(c) perfectly agree, showing zero back-scattering, thus further verifying the validity of our theory.